OpenStudy (anonymous):

suppose that, based on a sample, the 68% confidence interval for the mean of a population is (26, 32). what is the value of s/sqrtn? A. 5 B. 3 C. 2 D. 4

OpenStudy (anonymous):

\[\frac{ s }{ \sqrt{n} }\]

ganeshie8 (ganeshie8):

68% is 1 standard deviation away from mean

ganeshie8 (ganeshie8):

so, \(s\) = (32-26)/2 = 3

OpenStudy (anonymous):

so its B?

ganeshie8 (ganeshie8):

the question is asking u to find \(\large \frac{s}{\sqrt{n}}\),

ganeshie8 (ganeshie8):

you got \(\large s = 3\)

ganeshie8 (ganeshie8):

im not sure how to find \(n\)

OpenStudy (anonymous):

well 3 over something is closer than i was before so thank you

ganeshie8 (ganeshie8):

on second thoughts, \(\large \frac{s}{\sqrt{n}} = 3\) is right

ganeshie8 (ganeshie8):

correct option is B

ganeshie8 (ganeshie8):

\(3\) is the margin of error for given confidence interval, its not standard deviation

ganeshie8 (ganeshie8):

and, margin of error for 68% data is simply \(\large \frac{s}{\sqrt{n}}\)

OpenStudy (anonymous):

ahaha thank you, it was 3, i passed the test

ganeshie8 (ganeshie8):

yayyy! wonderful xD congrats !!